# UAV See, UGV Do: Aerial Imagery and Virtual Teach Enabling Zero-Shot Ground Vehicle Repeat

### v1

@@ Abstract
This paper presents Virtual Teach and Repeat (VirT&R): an extension of the Teach and Repeat (T&R) framework that enables GPS-denied, zero-shot autonomous ground vehicle navigation in untraversed environments. VirT&R leverages aerial imagery captured for a target environment to train a Neural Radiance Field (NeRF) model so that dense point clouds and photo-textured meshes can be extracted. The NeRF mesh is used to create a high-fidelity simulation of the environment for piloting an unmanned ground vehicle (UGV) to virtually define a desired path. The mission can then be executed in the actual target environment by using NeRF-derived point cloud submaps associated along the path and an existing LiDAR Teach and Repeat (LT&R) framework. We benchmark the repeatability of VirT&R on over 12 km of autonomous driving data using physical markings that allow a sim-to-real lateral path-tracking error to be obtained and compared with LT&R. VirT&R achieved measured root mean squared errors (RMSE) of 19.5 cm and 18.4 cm in two different environments, which are slightly less than one tire width (24 cm) on the robot used for testing, and respective maximum errors were 39.4 cm and 47.6 cm. This was done using only the NeRF-derived teach map, demonstrating that VirT&R has similar closed-loop path-tracking performance to LT&R but does not require a human to manually teach the path to the UGV in the actual environment.
@@ I INTRODUCTION
The proposed Virtual Teach and Repeat (VirT&R) system is implemented, tested, and evaluated on a Clearpath Warthog UGV. Autonomous repeats spanning over 12 km were conducted across a variety of environments, where Fig. 1 illustrates the process for creating a virtual teach map and then using a Clearpath Warthog to execute a mission without previously entering the environment. Path-tracking performance is then assessed and compared with an existing LT&R approach used previously [5]. To the best of the authors’ knowledge, this work presents the first demonstration of zero-shot closed-loop driving in outdoor environments by combining volumetric rendering of aerial imagery and a LiDAR sensor.
@@ I-A1 Teach and Repeat
The T&R navigation algorithm includes a teaching phase where local submaps are constructed through manual driving and followed in an autonomous repeat phase. The pose graph framework of T&R uses any sensor capable of providing a geometric world representation for localization [1, 2, 3, 4, 17, 18], and in this paper, we exploit this capacity to make use of the NeRF point cloud. By constructing a topometric pose graph, T&R retains the ability to achieve metric localization to the path without global consistency, which has been demonstrated on a UAV with vision sensors as well [7, 6]. However, the only attempts made towards cross-modal T&R have been with Radar and LiDAR sensors [19, 3], and these still required a manual teach phase in the target environment.
@@ II Methodology
The repeating phases of VirT&R and LT&R are identical, as this work presents an expansion to the algorithm [3] by enabling a new way to teach. Initial scene capture consists of flying a small UAV at various altitudes to gather different viewpoints of a scene for NeRF training data. A detailed mesh of the environment produced from this NeRF is used to create a simulator for driving the UGV to obtain the desired path through the environment as relative transformations. Subsequently, a dense point cloud from the NeRF is used to construct a sequence of small submaps which become associated to vertices along the desired path. Our method of teaching paths for the UGV enables multiple new paths to be virtually piloted through the captured scene without the UGV entering the environment beforehand. The capability of teaching multiple paths for the UGV to execute, given only one instance of a UAV capturing the scene, is a key benefit of our method.
@@ II-A Virtual Teach Pipeline
The VirT&R pipeline, illustrated in Fig. 3, outlines the key steps involved in enabling a UGV to navigate autonomously in an environment it has never physically traversed. The VirT&R pipeline consists of three main phases: a data capture and offline pre-processing phase to obtain imagery and reconstruct a scene, a simulated pilot driving phase to create the desired path, and an online execution phase where the robot completes a real mission. The offline VirT&R map creation method is as follows:
@@ II-A1 Scene Construction
The offline phase begins with UAV-based data collection, where aerial imagery is captured with a view looking straight down from approximately 20 m, looking forward at approximately 5-10 m, and looking at a 45° angle inwards on the scene at approximately 15 m. These images serve as the input for generating a 3D model of the environment and can optionally be post-processed with associated GPS poses so the NeRF model retains an exact real-world scale. However, this step is not necessary for the pipeline, as recovering a scaled scene can be done with methods such as the ICP algorithm. The reconstruction and optional GPS scaling are done using Colmap [41], which creates the inputs for the NeRF model that correspond to the training images and took 43 minutes for the largest scene with 1555 960x540 images. The four modeled scenes range from approximately 80 x 80 m to 150 x 100 m.
@@ II-A2 NeRF Model Training
The training is done on an NVIDIA GeForce RTX 3090 GPU, and took 22 minutes for the largest scene with 1555 960x540 images. We obtain two key representations from the trained NeRF model: a dense point cloud and a textured mesh that can be seen in Fig. 2 for the UTP Parking Lot Loop. The resulting representation of the environment is essential for subsequent localization and navigation tasks as the point cloud provides a rich, LiDAR-like representation of the environment, while the mesh facilitates visually accurate, simulation-based trajectory planning.
The mesh is imported into Gazebo [43] after being extracted from the NeRF model with the point cloud. The virtual teach pass can then be driven using a joystick controller and extracted as the absolute starting position and subsequent relative transformations in the Gazebo world frame. This simulated environment allows an operator to define a trajectory using a physics-based model of the Clearpath Warthog, replicating real-world driving and terrain assessment.
@@ II-C Evaluation and Ground Truth
The method for evaluating accuracy and path-tracking performance in VirT&R differs from prior work [1, 17, 18, 2, 3] because the teach phase is conducted entirely in a virtual environment, meaning we do not have access to GPS data for the teach path, which is commonly used to quantify the localization error. Thus, traditional validation methods that compare GPS trajectories of the teach and repeat phases are not applicable in this case.
To establish reference points for assessing repeat accuracy, an LT&R teach pass was conducted during which the robot was driven along a route and periodically stopped to spray paint marks near the rear-right tire. These marks were then captured in the NeRF model when training it from UAV imagery, allowing them to appear as features on the generated mesh. During simulation, the robot was piloted alongside these markings as closely to the distance they were painted from the wheel as possible to generate a comparable VirT&R teach pass. Fig. 4 illustrates the jig used to apply and measure the marks during the real repeat.
@@
To further validate performance, we repeated to the LT&R and VirT&R teach maps respectively, so that mark offset measurements could be taken for both algorithms, enabling a direct comparison between the two methods. Although aligning the robot to the marks in NeRF is inherently imperfect, this challenge mirrors the real-world difficulty of precisely following a physical reference path. To account for human error in the pilot simulation, a ruler model was attached to the Warthog in Gazebo, allowing measurements from the Warthog’s tire to the spray paint mark to be taken in simulation as well. By factoring in the sim-to-real measurement differences for the spray paint marks, we can better quantify the lateral path-tracking error by subtracting the error attributable to pilot imprecision from the VirT&R measurements, thereby providing a more accurate benchmark of performance and the sim-to-real gap between virtual path definition and the desired real-world mission.
@@ III-A Experimental Platforms
A Clearpath Warthog UGV [14], shown in Fig. 1, is equipped with an Ouster OS1-128 LiDAR [16] and a NovAtel SMART6 RTK-GPS system [45]. The Warthog is used to run LT&R on the teach maps generated from NeRF results. The DJI Phantom 4 Pro [15] is used for aerial imaging and uses the standard 20 MP CMOS sensor camera with an 84° field of view. Images for NeRF training were taken at 2.5 Hz (approximately every 0.5 m on average), and relative transformations from driving on the NeRF mesh were recorded at 20 Hz.
III-B Evaluation Routes
VirT&R was tested on the four routes in Fig. 5 at the University of Toronto Institute for Aerospace Studies campus. The UTIAS Parking Loop and UTP Loops consist of entirely paved paths, buildings, and cars. The Mars Dome route consists of a partially grass/snow-covered path, slight elevation changes, more vegetation, and fencing. Detailed visual cues embedded in the NeRF models of these environments include salt stains, tire tracks, puddles, and shadows, which allow for high-fidelity terrain assessment to be done by the pilot in the Gazebo simulation.
The Mars Dome was a particularly difficult environment due to 0.5 m of snow build-up, which required a path to be shoveled for the Warthog to be able to traverse part of the environment. This resulted in some parts of the route having less than 30 cm of lateral clearance during tight turns. Additionally, in the UTP Loop, two routes were taught and repeated to showcase one of the benefits of the virtual teach pass: multiple routes through one environment without any new scene data capture. Table I presents the environments and corresponding distances driven.
@@ IV Results
Qualitative assessment of repeat performance on all routes reveals that the robot can reliably follow the virtual path over long distances, navigate narrow openings, complete tight turns, and consistently hit specific marks. The NeRF meshes provide sufficient detail to infer scene characteristics—including potential dynamic elements—and the point clouds accurately represent the environment, even in predominantly flat areas. The UGV follows a smooth route and is resilient to both stationary and moving cars, pedestrians, and melted snow banks.
@@ IV-A LiDAR Teach and Repeat Baseline Comparison
To quantitatively compare VirT&R to LT&R, we implemented our physical marking-based evaluation methodology to obtain the lateral path-tracking RMSE, since GPS data is not available for the virtual teach phase. Measurements for the LiDAR and VirT&R repeats were taken five times each for the Mars Dome and UTIAS Parking Lot Loops, using identical controller parameters. Due to the nature of physically measuring these marks, an uncertainty of $\pm$4 cm has been associated with the readings to account for the measuring tape’s precision as well as the human error in placing the jig and stopping the Warthog next to the mark. Thus, we expect the LT&R repeats to agree with recently published [5] data within this uncertainty envelope.
@@
The spray paint mark offset measurements taken during VirT&R and LT&R repeats for the Mars Dome and UTIAS Parking Loops can be seen in Fig. 6, where the Mars Dome Loop had eight marks, and the UTIAS Parking Loop had twelve marks. Although the offset is consistently greater for VirT&R, the spread of the measurements is less than 10 cm for all marks, suggesting that VirT&R’s localization performs similarly to LT&R, allowing for the robot to repeat in its tracks—a finding that was observed qualitatively throughout testing, as faint tracks were visible after going through snow and puddles.
A summary of the lateral RMSE and maximum error values for VirT&R and LT&R, determined through physical measurement, can be seen in Table II, which also lists the T&R-estimated lateral and maximum errors obtained from the signed distances between every vertex in the teach and repeat pose graphs. The difference between the LT&R estimated and measured values is within the uncertainty associated with this method and indicates that physically measuring lateral path-tracking error is a reliable assessment. Additionally, LT&R estimated performance also aligns with a previous benchmark [5]. The measured lateral RMSE for VirT&R was 18.4 cm in the UTIAS Parking Loop and 19.5 cm in the Mars Dome Loop, which both fall within the range of LT&R maximum errors. The estimated VirT&R error for the Mars Dome Loop is likely higher than the measured value because the internal T&R estimate considers all vertices in the pose graph of a repeat, and the spray paint mark was likely not at the specific vertex with the highest offset, particularly as it was intentionally driven close to. The maximum VirT&R measured lateral path-tracking errors are 47.6 and 39.4 for the UTIAS Parking Loop and Mars Dome Loop respectively, signifying there is less accuracy with VirT&R than with LT&R, as expected. Overall, these results highlight that localization via NeRF-generated point clouds is a comparable method.
Figure 7: The repeatability of VirT&R is assessed by comparing two different repeats made using the NeRF-based virtual teach pose graphs for each of the four routes. The relative lateral path-tracking error between these two repeat trials is highlighted using colors ranging from purple (small error) to yellow (large error). The lateral error is observed to increase during tight turns or quick, precise maneuvering.
@@ IV-B VirT&R Repeatability
The repeatability of VirT&R is assessed by comparing two different VirT&R repeats for the Mars Dome, UTIAS Parking Lot, and both UTP Loops. In lieu of GPS data, signed distances between the vertices in the pose graph chains of respective repeats were used to determine the relative deviation between repeats to NeRF-generated teach maps. Given this same internally estimated metric was within the uncertainty of the measured lateral RMSE for the Mars Dome and UTIAS Parking Lot Loops, it is deemed to be a useful assessment.
Across the four routes seen in Fig. 7, the lateral error between two repeats increased from nearly zero to 55.4 cm, with the UTIAS Parking Loop exhibiting the best repeatability and the Mars Dome Loop having the worst. The portion of the Mars Dome Loop where the error is higher is also the area of the path with trees and slightly less structure. Additionally, during the Mars Dome virtual teach, intentionally rapid heading changes were driven in this area as a performance test.
The maximum relative error for all routes driven is found in this swerving section of the Mars Dome, near spray paint mark #4, which has the highest measured absolute lateral path-tracking error for this loop of 39.4 cm, seen in Table II. The largest errors consistently occur along sharp or more frequent turns on the routes, suggesting that the NeRF-based localization layer may not currently be robust enough to handle these maneuvers reliably. Overall, it can be seen that the lateral error between VirT&R repeat trials is mostly small along the path and increases slightly during tight turns or quick maneuvers.

### v2

@@ Abstract
This paper presents Virtual Teach and Repeat (VirT&R): an extension of the Teach and Repeat (T&R) framework that enables GPS-denied, zero-shot autonomous ground vehicle navigation in untraversed environments. VirT&R leverages aerial imagery captured for a target environment to train a Neural Radiance Field (NeRF) model so that dense point clouds and photo-textured meshes can be extracted. The NeRF mesh is used to create a high-fidelity simulation of the environment for piloting an unmanned ground vehicle (UGV) to virtually define a desired path. The mission can then be executed in the actual target environment by using NeRF-generated point cloud submaps associated along the path and an existing LiDAR Teach and Repeat (LT&R) framework. We benchmark the repeatability of VirT&R on over 12 km of autonomous driving data using physical markings that allow a sim-to-real lateral path-tracking error to be obtained and compared with LT&R. VirT&R achieved measured root mean squared errors (RMSE) of 19.5 cm and 18.4 cm in two different environments, which are slightly less than one tire width (24 cm) on the robot used for testing, and respective maximum errors were 39.4 cm and 47.6 cm. This was done using only the NeRF-derived teach map, demonstrating that VirT&R has similar closed-loop path-tracking performance to LT&R but does not require a human to manually teach the path to the UGV in the actual environment.
@@ I INTRODUCTION
The proposed Virtual Teach and Repeat (VirT&R) system is implemented, tested, and evaluated on a Clearpath Warthog UGV. Autonomous repeats spanning over 12 km were conducted across a variety of environments, where Fig. 1 illustrates the process for creating a virtual teach map and using a Clearpath Warthog to execute a mission without previously entering the environment. Path-tracking performance is then assessed and compared with an existing LT&R approach used previously [5]. To the best of the authors’ knowledge, this work presents the first demonstration of zero-shot closed-loop driving in outdoor environments by combining volumetric rendering of aerial imagery and a LiDAR sensor.
@@ I-A1 Teach and Repeat
The T&R navigation algorithm includes a teaching phase where local submaps are constructed through manual driving and followed in an autonomous repeat phase. The pose graph framework of T&R uses any sensor capable of providing a geometric world representation for localization [1, 2, 3, 4, 17, 18], and in this paper, we exploit this capacity to make use of the NeRF-generated point cloud. By constructing a topometric pose graph, T&R retains the ability to achieve metric localization to the path without global consistency, which has been demonstrated on a UAV with vision sensors as well [7, 6]. However, the only attempts made towards cross-modal T&R have been with Radar and LiDAR sensors [19, 3], and these still required a manual teach phase in the target environment.
@@ II Methodology
The online repeating phases of VirT&R and LT&R are identical as we present an expansion to the algorithm [5] by enabling a new way to teach, and existing downstream modules of T&R that are run live during a repeat handle motion planning and control. Initial scene capture consists of flying a small UAV at various altitudes to gather different viewpoints of a scene for NeRF training data. A detailed mesh of the environment produced from this NeRF model is used to create a simulator for driving the UGV to obtain the desired path through the environment as relative transformations. Subsequently, a dense point cloud from the NeRF model is used to construct a sequence of small submaps, which become associated with vertices along the desired path. Our method of teaching paths for the UGV enables multiple new paths to be virtually piloted through the captured scene without the UGV entering the environment beforehand. The capability of teaching multiple paths for the UGV to execute, given only one instance of a UAV capturing the scene, is a key benefit of our method.
@@ II-A Virtual Teach Pipeline
The VirT&R pipeline, illustrated in Fig. 3, outlines the key steps involved in enabling a UGV to navigate autonomously in an environment it has never physically traversed. The VirT&R pipeline consists of three main phases: a data capture and offline pre-processing phase to obtain a scene reconstruction and localization layer, a simulated pilot driving phase to create the desired path, and an online execution phase where the robot completes a real mission. The offline VirT&R map creation method is as follows:
@@ II-A1 Scene Construction
The offline phase begins with UAV-based data collection, where aerial imagery is captured with a view looking straight down from approximately 20 m, looking forward at approximately 5-10 m, and looking at a 45° angle inwards on the scene at approximately 15 m. These images serve as the input for generating a 3D model of the environment and can optionally be post-processed with associated GPS poses so the NeRF model retains an exact real-world scale. However, this step is not necessary for the pipeline, as recovering a scaled scene can be done with methods such as the ICP algorithm. The reconstruction and optional GPS scaling are done using Colmap [41], which creates the inputs for the NeRF model that correspond to the training images and took 43 minutes for the largest scene with 1555 960x540 images. The modeled scenes range from approximately 80 x 80 m to 150 x 100 m.
@@ II-A2 NeRF Model Training
The training is done offline using UAV imagery on an NVIDIA GeForce RTX 3090 GPU, and took 22 minutes for the largest scene with 1555 960x540 images. We obtain two key representations from the NeRF model once training has finished: a dense point cloud and a textured mesh that can be seen in Fig. 2 for the UTP Parking Lot Loop. The resulting representations of the environment are essential for subsequent localization and navigation tasks, as the point cloud provides a rich LiDAR-like representation of the environment, while the mesh facilitates visually accurate simulation-based trajectory planning.
The mesh is imported into Gazebo [43] after being extracted from the NeRF model with the point cloud. The virtual teach pass can then be driven using a joystick controller and saved as the absolute starting position and subsequent relative transformations in the Gazebo world frame. This simulated environment allows an operator to define a trajectory using a physics-based model of the Clearpath Warthog, replicating real-world driving and terrain assessment.
@@ II-C Evaluation and Ground Truth
The method for evaluating accuracy and path-tracking performance in VirT&R differs from prior work [1, 17, 18, 2, 3] because the teach phase is conducted entirely in a virtual environment. As such, we do not have the GPS data for teach paths commonly used to quantify path-tracking error (PTE).
In an attempt to produce ‘pseudo-GPS’ data for evaluation, we recovered the path taught with the Gazebo simulation in an Earth Centered, Earth Fixed (ECEF) coordinate frame using the GPS data from the UAV that gave the accurate scale to the NeRF model. GPS data for repeats to the virtual teach map was also recorded using a NovAtel OEM7 GPS [45] on all routes. Analysis of the resulting GPS-measured PTE revealed drastically elevated values that did not align with our qualitative observations or the physical markings. It is likely that multiple sources of error, such as GPS receiver differences, a weeks-long gap between GPS measurements, Colmap and NeRF reconstruction inaccuracies, and the lack of GPS post-processing on UAV data, were compounded in creating the pseudo-GPS data. Thus, it was deemed impractical for use in measuring the actual PTE from control and localization.
The lack of GPS data for conventional evaluation led us to establish a physical marking-based sim-to-real path-tracking assessment method. Physical markings in the environment are the most suitable tool to assess the UGV’s ability to track the virtual teach path because they ensure the PTE accounts for our own errors in piloting the UGV in simulation, as well as localization and control errors, not any that may result from the scene reconstruction process. This is due to the markings appearing in the drone imagery, the Colmap reconstruction and NeRF model, and the photo-textured mesh, meaning any drift or distortion from scene reconstruction is applied equally to the markings, the virtual teach path driven alongside them, and the point cloud used for localization. A visualization of how reconstruction error is implicitly canceled out can be seen in Fig. 4, where the use of local, potentially drifted submaps made from the virtually-generated point cloud are localized to the real environment to enable the desired virtually taught path to be repeated.
Figure 4: A visualization of a desired real world path with spray paint markings, an exaggeration of what potential drift could do to the virtually-generated mesh and point cloud (and thus the virtual teach path), and the desired path being repeated due to equally applied drift so the correct teach path is recovered in the real world.
To implement this methodology, an LT&R teach pass was conducted on the Mars Dome and Parking Loops during which the robot was driven and periodically stopped to spray paint marks near the rear-right tire. These marks were then captured in the UAV footage, scene reconstruction, and mesh, so they could be virtually piloted alongside of as closely to the distance they were painted from the tire as possible, to generate a comparable VirT&R teach pass for offsets to be measured during a real repeat. The jig used to apply and measure the markings in the real world is shown in Fig. 5.
@@
To further validate performance, we repeated to the LT&R and VirT&R teach maps respectively, so that mark offset measurements could be taken for both algorithms, enabling a direct comparison between the two methods. Although aligning the UGV with the marks in the simulation is inherently imperfect, this challenge mirrors the real-world difficulty of precisely following a physical reference path. However, it does prevent us from comparing the exact same paths, as the virtually driven path is likely not exactly what the LT&R path used to place the markings is. To account for this human error, a ruler model was attached to the Warthog in Gazebo, allowing measurements from the Warthog’s tire to the spray paint mark to be taken in simulation as well. We can better quantify the PTE by subtracting the error attributable to pilot imprecision from the VirT&R repeat offset measurements, thereby providing a more accurate benchmark of performance and the sim-to-real gap.
@@ III-A Experimental Platforms
A Clearpath Warthog UGV [14], shown in Fig. 1, is equipped with an Ouster OS1-128 LiDAR [16] and a NovAtel SMART6 RTK-GPS system [45]. The Warthog is used to run LT&R on the teach maps generated from NeRF results. The DJI Phantom 4 Pro [15] is used for aerial imaging and uses the standard 20 MP CMOS sensor camera with an 84° field of view. Images for NeRF training were taken at 2.5 Hz (approximately every 0.5 m), and relative transformations from virtual driving on the mesh were recorded at 20 Hz.
@@
III-B Evaluation Routes
VirT&R was tested on the four routes in Fig. 6 at the University of Toronto Institute for Aerospace Studies (UTIAS) campus. The UTIAS Parking Loop and UTP Loops consist of entirely paved paths, buildings, and cars. The Mars Dome route consists of a partially grass/snow-covered path, slight elevation changes, more vegetation, and fencing. Detailed visual cues embedded in the NeRF models of these environments include salt stains, tire tracks, puddles, and shadows, which enable high-fidelity terrain assessment to be done by the pilot in simulation.
The Mars Dome was a particularly difficult environment due to 0.5 m of snow build-up, which required a path to be shoveled for the Warthog to be able to traverse part of the environment. This resulted in some parts of the route having less than 30 cm of lateral clearance during tight turns. Additionally, in the UTP Loop, two routes were taught to showcase one of the benefits of the VirT&R: teaching multiple routes through one environment without any new scene data capture. Table I presents the environments and corresponding distances driven.
@@ IV Results
Qualitative assessment of repeat performance on all routes reveals that the robot can reliably follow the virtual path over long distances, navigate narrow openings, complete tight turns, and consistently hit specific marks. The NeRF meshes provide sufficient detail to infer scene characteristics—including potential dynamic elements—and the point clouds accurately represent the environment, even in predominantly flat areas. The UGV follows a route smoothly and is resilient to both stationary and moving cars, pedestrians, and melted snow banks while repeating with a virtual teach map.
@@ IV-A LiDAR Teach and Repeat Baseline Comparison
To quantitatively compare VirT&R to LT&R, we implemented our physical marking-based evaluation methodology to obtain the lateral path-tracking RMSE. Measurements for the LiDAR and VirT&R repeats were taken five times each for the Mars Dome and UTIAS Parking Lot Loops, using identical controller parameters. Due to the nature of physically measuring these marks, an uncertainty of $\pm$4 cm has been associated with the readings to account for the measuring tape’s precision as well as human error when placing the jig and stopping the Warthog next to the mark. Thus, we expect the LT&R repeats to agree with recently published [5] data within this uncertainty envelope.
@@
The marking offset measurements taken during VirT&R and LT&R repeats for the Mars Dome and UTIAS Parking Loops can be seen in Fig. 7, where the Mars Dome Loop had eight marks, and the UTIAS Parking Loop had twelve. Although the offset is consistently greater for VirT&R, the spread of the measurements is less than 10 cm for all marks, suggesting that VirT&R’s localization performs similarly to LT&R, allowing for the UGV to repeat in its tracks—a finding we observed qualitatively throughout testing, as faint tracks were visible after going through snow and puddles.
A summary of the lateral RMSE and maximum error values for VirT&R and LT&R, determined through physical measurement, can be seen in Table II, which also lists the T&R-estimated lateral and maximum errors obtained from the signed distances between every vertex in the teach and repeat pose graphs. The difference between the LT&R estimated and measured values is within the uncertainty associated with this method and indicates that physically measuring lateral path-tracking error is a reliable assessment. Additionally, LT&R estimated performance also aligns with a previous benchmark [5]. The measured lateral RMSE for VirT&R was 18.4 cm in the UTIAS Parking Loop and 19.5 cm in the Mars Dome Loop, which both fall within the range of LT&R maximum errors. The estimated VirT&R error for the Mars Dome Loop is likely higher than the measured value because the internal T&R estimate considers all vertices in the pose graph of a repeat, and the spray paint mark was likely not at the specific vertex with the highest offset, particularly as it was intentionally driven close to. The maximum VirT&R measured lateral path-tracking errors are 47.6 and 39.4 cm for the UTIAS Parking Loop and Mars Dome Loop respectively, signifying there is less accuracy with VirT&R than with LT&R, as expected. Overall, these results highlight that localization via NeRF-generated point clouds is a comparable method.
@@ IV-B VirT&R Repeatability
The repeatability of VirT&R is assessed by comparing two repeats each for the Mars Dome, UTIAS Parking Lot, and UTP Loops. This comparison is useful to gauge the affect that non-stationary objects and environmental changes have over time on the executed repeat. In lieu of GPS data, signed distances between vertices in the pose graph chains of respective repeats were used to determine the relative deviation between repeats to virtual teach maps. Given this same internally estimated metric was within the uncertainty of the measured lateral RMSE for the Mars Dome and UTIAS Parking Lot Loops, it is deemed to be a useful assessment.
Figure 8: VirT&R repeatability is assessed by comparing two repeats made using virtual teach maps for each of the four routes. The relative path-tracking error between repeats is highlighted using colors ranging from purple (small error) to yellow (large error). Lateral error is observed to increase during tight turns or quick, precise maneuvering.
Across the four routes in Fig. 8, the lateral error between two repeats increased from nearly zero to 55.4 cm, with the UTIAS Parking Loop exhibiting the best repeatability and the Mars Dome Loop having the worst. The portion of the Mars Dome Loop where the error is higher is also the area of the path with trees and slightly less structure. Additionally, during the virtual teach, intentionally rapid heading changes were driven here as a performance test.
The maximum relative error for all routes is found in this swerving section of the Mars Dome Loop, near marking #4, which has the highest measured absolute lateral path-tracking error for this loop of 39.4 cm, seen in Table II. The largest errors consistently occur along sharp or more frequent turns, suggesting that the NeRF-generated localization layer may not currently be robust enough to handle these maneuvers reliably. Overall, it can be seen that the lateral error between VirT&R repeats is mostly small along the path and increases during tight turns or quick maneuvers.